OpenStudy (alina123):

Really easy, just check my answer. If f (x) = 4x^2+3 and g (x ) = 1/3x+4 , find ( f of g^-1)(3). I got 39 @rishavraj

rishavraj (rishavraj):

did u calculate the \[g^{-1}(x)\]

OpenStudy (alina123):

Yes, and i got 3x-12 = y

rishavraj (rishavraj):

hold on solving :P

OpenStudy (alina123):

did u solve?

OpenStudy (color):

lets say g(x)=y \(g^{-1}(x)=\large\frac{1}{3(y-4)}\)

OpenStudy (alina123):

um ok?

rishavraj (rishavraj):

u got options...coz i am getting fractions

OpenStudy (alina123):

ok to find g^-1 you simply find the inverse of g (x)

OpenStudy (color):

ok now just plug in \(g^{-1}(x)\) into f(x) in place of x to get (f of g^-1)(x)

OpenStudy (alina123):

so thats 3x-12, then plug in 3, then use that value in f

OpenStudy (alina123):

wait im confused

rishavraj (rishavraj):

@Alina123 there???

rishavraj (rishavraj):

whts the value of g-1(x) ??

OpenStudy (alina123):

its just the inverse of g(x)

rishavraj (rishavraj):

lol i dint mean tht.... u calculated g-1(x) right??/

OpenStudy (alina123):

yeh i got 3x-12

OpenStudy (alina123):

Hello

rishavraj (rishavraj):

yeah

OpenStudy (alina123):

um please help me with my question i dont have time

rishavraj (rishavraj):

\[g^{-1} (x) = \frac{ 1 - 4x }{ 3x }\]

OpenStudy (alina123):

this doesnt help what will my answer overall be?

rishavraj (rishavraj):

no no .... now u need to calculate \[g^{-1}(3)\]

OpenStudy (alina123):

ok what do u get then

rishavraj (rishavraj):

787/81

OpenStudy (alina123):

9.7

rishavraj (rishavraj):

keep it in fraction

OpenStudy (shadowlegendx):

Inverse the g function, input it into the function of x, with x being 3 as per g(3)

OpenStudy (shadowlegendx):

Do you know how to inverse a function?

OpenStudy (alina123):

yes

OpenStudy (alina123):

i get -3 then

OpenStudy (shadowlegendx):

Let me try solve it

OpenStudy (shadowlegendx):

Inverse of g(x) \[g (x )^-1 = \frac{ 1 }{ 3 }x +4 \] \[y = \frac{ 1 }{ 3 }x + 4\] Switch every x with every y \[x = \frac{ 1 }{ 3 }y + 4 \] Solve for y \[y = 3(x -4)\] Therefore, \[g(x)^{-1} = 3(x-4)\]

OpenStudy (shadowlegendx):

Does this make sense to you?

OpenStudy (alina123):

yes i get this part

OpenStudy (alina123):

i then substitute 3 and get -3

rishavraj (rishavraj):

@Alina123 it was \[g(x) = \frac{ 1 }{ 3 }x + 4\] not \[g(x) = \frac{ 1 }{ 3x + 4 }\] -_-

OpenStudy (alina123):

No, cause @shadowlegendx his way was the way i was going for

OpenStudy (shadowlegendx):

I got 39 as well. Inputting -3 from g(3)^-1